### v1

# OBLIVIATE: Robust and Practical Machine Unlearning for Large Language Models

Abstract
Large language models (LLMs) trained over extensive corpora risk memorizing sensitive, copyrighted, or toxic content. To address this, we propose OBLIVIATE, a robust unlearning framework that removes targeted data while preserving model utility. The framework follows a structured process: extracting target tokens, building retain sets, and fine-tuning with a tailored loss function comprising three components—masking, distillation, and world fact. Using low-rank adapters (LoRA), it ensures efficiency without compromising unlearning quality. We conduct experiments on multiple datasets, including the Harry Potter series, WMDP, and TOFU, using a comprehensive suite of metrics: forget quality (new document-level memorization score), model utility, and fluency. Results demonstrate its effectiveness in resisting membership inference attacks, minimizing the impact on retained data, and maintaining robustness across diverse scenarios. [^1] Our code is available at https://anonymous.4open.science/r/OBLIVIATE_unlearning_LLM-0C84.
OBLIVIATE: Robust and Practical Machine Unlearning
for Large Language Models
Xiaoyu Xu1, Minxin Du1, Qingqing Ye1, Haibo Hu1, 1The Hong Kong Polytechnic University, xiaoyu0910.xu@connect.polyu.hk, {minxin.du, qqing.ye, haibo.hu}@polyu.edu.hk
1 Introduction
Figure 1: Overview of OBLIVIATE, a robust and practical unlearning framework for LLMs
The rapid expansion of training data for large language models (LLMs) has driven significant advancements across various domains. However, the tendency of LLMs to memorize training corpora raises critical ethical and security concerns, such as generating sensitive, harmful, or copyrighted content Nasr et al. (2023); Karamolegkou et al. (2023); Wen et al. (2023). These issues highlight the need to adapt LLMs to diverse security environments and meet user and industry-specific requirements, with regulations like the EU’s Right to be Forgotten Ginart et al. (2019) further emphasizing their importance. In response, machine unlearning has emerged as a promising solution to mitigate these risks Yao et al. (2024); Jang et al. (2023); Eldan and Russinovich (2023); Pawelczyk et al. (2024); Li et al. (2024b, a, 2025). Unlearning ensures that models behave as if specific data were never included in the training sets Bourtoule et al. (2021), effectively reducing sensitive information leakage and aligning LLMs with legal standards.
Current LLM unlearning methods generally fall into three categories: fine-tuning Yao et al. (2024), prompt-based Liu et al. (2024a), and task arithmetic Ilharco et al. (2023); Ji et al. (2024). Fine-tuning-based methods update model parameters to maximize the unlearning effect while maintaining performance on retained data. In contrast, prompt-based and task arithmetic methods modify input prompts or output logits without altering the model’s parameters. Among these, fine-tuning-based methods often achieve superior results.
However, common fine-tuning approaches, such as gradient ascent (GA), random label fine-tuning, and adversarial sample-based methods Yao et al. (2024), face several limitations. First, Shi et al. (2024) shows that unlearned data can often be recovered via membership inference attacks (MIAs), suggesting that memorized information is not fully erased. Second, balancing effective unlearning with performance preservation on retained data remains challenging. Techniques like gradient descent or KL-divergence on retained data often fail to maintain model utility in real-world scenarios, particularly due to the impracticality of defining clear retain set boundaries without access to proprietary training corpora. Finally, existing evaluations are often insufficient, lacking comprehensiveness and reliability in verifying whether the forget set has been fully removed and whether model performance remains intact Liu et al. (2024b).
To address the challenges, we propose OBLIVIATE, a robust and practical LLM unlearning framework that effectively removes target data while preserving model performance (e.g., on various downstream tasks) and fluency–defined as the ability to generate coherent and precise responses–on the retain set. Figure 1 outlines OBLIVIATE with three critical loss functions: masked loss for the forget set, and distillation and world fact losses for the retain set. Additionally, we utilize low-rank adapters (LoRA) Hu et al. (2022) for fine-tuning efficiency.
To optimize forget quality for strict regulatory compliance Ginart et al. (2019), we introduce a masked loss that enforces zero-generation probability for targeted content, facilitating “aggressive” forgetting, inspired by multimodal unlearning Li et al. (2024a). However, this aggressive approach can degrade model performance and fluency on the retain set, often producing incoherent outputs Thaker et al. (2024). To mitigate potential catastrophic forgetting, we incorporate two additional losses: distillation and world fact. The distillation loss aligns the model with teacher models trained on related documents, preserving performance and fluency on the retain set. The world fact loss uses encyclopedic data (e.g., WikiText Merity et al. (2017)) to maintain general factual knowledge. These two extra losses allow the model to perform context-aware unlearning—selectively forgetting sensitive information in harmful contexts while preserving knowledge in benign contexts and triggering forgetting only when necessary. As shown in Table 15, they can prevent indiscriminate erasure of unrelated knowledge.
We validate the robustness and effectiveness of OBLIVIATE across multiple datasets, demonstrating strong unlearning performance while maintaining model utility and fluency. To ensure comprehensive and reliable evaluation, we introduce a suite of metrics for assessing forget quality, model utility, and fluency. Our main contributions are:
I) We propose OBLIVIATE, an LLM unlearning framework that can effectively eliminate the influence of unlearning data while preserving the model’s performance and fluency on the retain set.
II) We introduce a masked loss that completely suppresses the generation of unlearning data, showing competitive effectiveness compared to other fine-tuning-based methods Yao et al. (2024).
III) To counteract the masked loss’s negative impacts, we devise distillation and world fact losses to preserve generic knowledge and model fluency.
IV) We conduct experiments on multiple datasets (with varying scope) to validate the performance of OBLIVIATE using a comprehensive evaluation suite of forget quality, model utility, and fluency.
2 Problem Formulation
Let $\mathcal{D}$ be a large training corpus, and let $\mathcal{D}_{f}\subseteq\mathcal{D}$ be the forget set to be unlearned, containing a set of $M$ documents $\{d_{i}\}_{i=1}^{M}$ (e.g., book, personal records). Each $d_{i}=\{x_{j}\}_{j=1}^{N}$ is a sequence of $N$ tokens. Given a model $\mathcal{M}$ trained on $\mathcal{D}$ using an algorithm $\mathcal{A}$, an unlearning algorithm $\mathcal{U}$ is applied to $\mathcal{M}$, with each $d_{i}$ as input, to produce an unlearned model $\mathcal{M}^{\prime}$, effectively removing the effects of $\mathcal{D}_{f}$.
Inspired by differential privacy Gupta et al. (2021); Sekhari et al. (2021); Neel et al. (2021); Du et al. (2023), the NeurIPS 2023 machine unlearning challenge [^2] https://unlearning-challenge.github.io/assets/data/Machine_Unlearning_Metric.pdf parameterizes unlearning by $(\epsilon,\delta)$, quantifying the difference between the distributions of $\mathcal{U}(\mathcal{M})$ and $\mathcal{A}(\mathcal{D}\setminus\mathcal{D}_{f})$. When $\epsilon=\delta=0$, $\mathcal{U}$ is exact unlearning—the output distributions are identical. While retraining achieves exact unlearning, it is computationally prohibitive for LLMs Luccioni et al. (2023); Zhang et al. (2023). For small, positive $\epsilon$ and $\delta$, $\mathcal{U}$ is approximate unlearning, offering a practical solution for real-world applications.
The theoretical framework is often not “applicable” to non-convex structures like LLMs Kim et al. (2021). Most current LLM unlearning studies rely on empirical evaluation rather than strict theoretical guarantees Eldan and Russinovich (2023); Maini et al. (2024); Li et al. (2024b); Gandikota et al. (2024). These evaluations typically compare the unlearned model to the retrained model on benchmark datasets (e.g., MMLU, MT-Bench), assessing metrics, such as forget quality and model utility Maini et al. (2024). We follow this evaluation strategy.
2.1 Scope of LLM Unlearning
LLM unlearning is motivated by three imperatives: copyright, privacy protection, and the mitigation of harmful outputs Liu et al. (2024b).
Copyright. To satisfy intellectual-property regulations, models must purge training data incorporated without authorization. Ongoing litigation involving OpenAI, Meta, and The New York Times underscores this need Small (2023). Experiments on the Harry Potter corpus show that targeted unlearning can remove copyrighted content and reduce legal exposure Eldan and Russinovich (2023).
Privacy. Unlearning curbs the memorization of personally identifiable information (PII) Jang et al. (2022); Carlini et al. (2023). The synthetic TOFU benchmark gauges how effectively private attributes can be removed Maini et al. (2024).
Harmful outputs. By erasing knowledge that allows toxic, discriminatory, or dangerous content, unlearning aligns the models with the values of society. Results on the WMDP dataset, which contains biorisk and cybersecurity material, demonstrate its efficacy Li et al. (2024b).
3 Methodology
3.1 Overview
We put forth OBLIVIATE, an LLM unlearning framework with two phases: i) pre-processing to identify target tokens for unlearning and create a retain set to preserve model performance and fluency (Section 3.2) and ii) fine-tuning using LoRA and a tailored unlearning loss (Section 3.3), which has three components: the masked loss suppressing the forget set $\mathcal{D}_{f}$ by enforcing zero-generation probabilities for target tokens, the distillation loss aiding in preserving model performance on the retain set by aligning the model with teacher models trained on related documents, and the world fact loss maintaining general factual knowledge by using encyclopedic sources like WikiText. To evaluate the “forget quality,” we introduce document-level memorization, a new metric to capture broader memorization behavior across documents.
3.2 Pre-processing
Identification of target (to-be-unlearned) tokens.
Li et al. (2024a) proposed two masking strategies for multimodal unlearning: token- and vocabulary-level, where the former selectively excludes specific tokens from loss computation, and the latter globally suppresses the probabilities of targeted concepts. To balance model behavior preservation with output suppression, the Dual Masked KL-divergence (DMK) loss was introduced, which applies both masking strategies during fine-tuning.
In contrast, we only employ vocabulary-level masking, implemented by zeroing out logits for target tokens before the softmax operation. After masking, we compute the masked target distribution using softmax and optimize a KL-divergence loss between the original and masked model outputs. This suppresses the target tokens’ probabilities without the need for token-level exclusion. Unlike Li et al. (2024a), we do not apply token-level masking due to its high costs for large-scale unlearning. It requires explicit identification of target tokens within individual sentences, which can be challenging and may disrupt semantic coherence.
For token identification, Li et al. (2024a) uses the next-token probability distribution, but our setting involves broader and more complex target concepts, such as entities, locations, events, and relationships in datasets (like the Harry Potter series). Enumerating all potential target tokens is impractical. Statistical methods, such as token frequency and probability Meeus et al. (2024), while efficient, often miss unique tokens. Named entity recognition (NER) Roy (2021) relies on predefined target sets to identify tokens. Instead, we leverage GPT-4o to “identify” target tokens through tailored prompts (see Appendix D). It combines the benefits of NER, such as prior knowledge and contextual understanding, with the scalability and efficiency of statistical approaches, allowing flexible token identification with minimal computational overhead. Based on the identified target tokens, we construct a masked loss for unlearning $\mathcal{D}_{f}$ in Section 3.3.
Remark. The use of GPT-4 for similar tasks has been explored in prior work Eldan and Russinovich (2023); Liu et al. (2024c); Shi et al. (2024); Maini et al. (2024). For example, Eldan and Russinovich (2023) employs GPT-4 to detect specific “anchored terms” while suggesting generic alternatives, and Shi et al. (2024) leverages GPT-4 to paraphrase sensitive answers for enhanced privacy.
The computational cost of GPT-4o-based token identification is also minimal, e.g., identifying target tokens across $400$ documents in the WMDP dataset takes $\sim$26s vs. $991$s for the entire unlearning process (Table 7). For longer or information-rich inputs, GPT-4o supports a context length of up to $128$k tokens [^3] https://platform.openai.com/docs/models/gpt-4o. Two strategies can process inputs exceeding this limit: i) splitting the document into multiple 128k-token segments, or ii) processing file-based inputs while incrementally parsing key tokens. We use the second one in our experiments.
While the GPT-4o-based approach may have limitations, particularly with subtle context-dependent expressions, our results show its effectiveness, outperforming several baselines. Further exploration of advanced approaches is left for future work.
Construct retain set.
We build a retain set with three document categories—generic, other-style, and world-fact—each containing $M$ documents, matching the size of the forget set $\mathcal{D}_{f}$.
Generic documents. To preserve performance on inputs resembling $\mathcal{D}_{f}$, we select full-length documents that mirror the semantics and token counts of each $d_{i}\in\mathcal{D}_{f}$. Candidates with the highest BM25 similarity, a probabilistic relevance metric Cheng et al. (2024), are chosen in Algorithm 1. A predefined retain set can substitute this selection step.
Other-style documents. These maintain domain competence while varying stylistic features. Consider the Harry-Potter series as a forget set, we add novels from distinct genres (e.g., historical or contemporary fiction). For non-narrative data, token-order shuffles of the generic documents suffice.
World-fact documents. When $\mathcal{D}_{f}$ includes general knowledge (e.g., geography, cuisine), we supplement the retain set with encyclopedic sources, such as WikiText Merity et al. (2017), to safeguard factual utility as in RMU Gandikota et al. (2024).
3.3 Tailored Unlearning Loss
The core of OBLIVIATE is a customized unlearning (or fine-tuning) loss function with three components, each targeting a specific document type.
Masked loss.
For input $d_{i}\in\mathcal{D}_{f}$, we set the probabilities of the target tokens in the output distribution to zero, resulting in a masked logits distribution. We introduce a masked loss using KL divergence to minimize the difference between the masked and original logits distributions.
Our approach prioritizes two objectives. First, we aim to optimize forget quality to meet strict regulatory compliance requirements Ginart et al. (2019), while maintaining (near-)optimal model utility and fluency. Second, previous studies Zhang et al. (2024b); Shi et al. (2024) show that “weaker” unlearning methods, such as NPO Zhang et al. (2024a) and WHP Eldan and Russinovich (2023), are more vulnerable to MIAs, highlighting the need for more aggressive forgetting mechanisms.
In contrast to the DMK loss Li et al. (2024a), which separately applies token- and vocabulary-level masking, ours directly zeros out the probabilities of target tokens, enforcing stricter alignment between the masked and original distributions. This results in a stronger, more focused forgetting effect, at the cost of increased aggressiveness. Applying DMK in our context would introduce significant computational overheads due to token-level masking and may unintentionally affect unrelated knowledge. Our approach mitigates these issues by employing a globally enforced masking strategy that is better suited for large-scale text unlearning.
Our masked loss is formulated as
$\mathcal{L}_{\text{Mk}}(P\|Q)=\sum_{d_{i}\in\mathcal{D}_{f}}P(\theta_{masked})\log\frac{P(\theta_{masked})}{Q(\theta)},$
where $P(\theta_{masked})$ and $Q(\theta)$ are the masked and original logits distributions, respectively.
The masked loss shares a similar goal of aggressive unlearning with GA Golatkar et al. (2020)–both utilize “negative” updates to remove unlearned data, which is essential for fully removing memorized knowledge in the forget set. To prevent catastrophic collapse or excessive unlearning, we introduce two auxiliary losses–distillation and world fact–as suggested by Yao et al. (2024). These two losses enable context-aware unlearning: selectively forgetting sensitive information in harmful contexts, preserving knowledge in benign contexts, and triggering forgetting only when necessary.
Figure 2: Token-level RMA vs. Our DRMA
Distillation loss.
To retain fluency after unlearning, we distill knowledge from two teacher models: one trained on generic documents, and the other on other-style data. For each forget-set example $x_{1}$ and its paired style counterpart $x_{2}$, we minimize the mean-squared error (MSE) between the student’s logits $P(\theta_{x_{1}})$ and the teachers’ logits $P^{\prime}(\theta_{x_{2}})$:
$\mathcal{L}_{\text{distillation}}=\mathbb{E}_{x_{1},x_{2}}\text{MSE}(P(\theta_{x_{1}}),P^{\prime}(\theta_{x_{2}})).$
MSE exploits the full soft distribution, offering smoother gradients and finer feature transfer than cross-entropy (CE) loss, which overweights the top class (see Appendix C). Aligning the student with both teachers suppresses over-frequent tokens (e.g., ‘a’) and sustains coherent, well-structured outputs.
World-fact loss.
Lastly, we apply an extra world-fact loss to preserve encyclopedic knowledge on inputs drawn from WikiText Merity et al. (2017). Specifically, we minimize the cross-entropy (CE) loss between the target model’s output distribution $P(\theta)$ and that of the original model $P^{\prime\prime}(\theta)$:
$\mathcal{L}_{\text{world fact}}=\mathbb{E}_{x\in\text{Wikipedia}}\text{CE}(P(\theta),P^{\prime\prime}(\theta)).$
Note that the CE loss fits such a categorical setting and follows precedent in factual-retention studies Gandikota et al. (2024); Gu et al. (2024). Aligning the two distributions can protect general-knowledge utility after unlearning.
Our final objective combines the three losses:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{forget}}+\lambda_{1}\mathcal{L}_{\text{distillation}}+\lambda_{2}\mathcal{L}_{\text{world fact}},$
where $\lambda_{1},\lambda_{2}$ are tunable hyperparameters. To unlearn $\mathcal{D}_{f}$, we thus apply LoRA to fine-tune the MLP and MHA layers of LLM using $\mathcal{L}_{\text{total}}$.
3.4 Document-level memorization
To directly capture the “memorization behavior” across $M$ documents (or sequences), we generalize the token-level Remnant memorization accuracy (RMA) Lee et al. (2024) to document level. Specifically, for $M$ documents, each with $n$ tokens, our document-level RMA (DRMA) is defined as
$\text{DRMA}=\frac{\sum_{i=1}^{M}\sum_{t=1}^{n-1}p_{\theta}(x_{t}\mid x_{<t})}{M},$
where $p_{\theta}(x_{t}\mid x_{<t})$ denotes the probability of outputting the $t$-th token $x_{t}$, conditioned on the preceding tokens $x_{<t}$ within a document: A lower DRMA value indicates reduced document memorization. It assesses broader patterns rather than individual tokens, offering a more holistic view (than RMA) of the ability to forget information, which is crucial for unlearning tasks involving multi-sequence contents; see Figure 2. Beyond DRMA, our experiments use other metrics (e.g., resistance to MIAs) to assess the forget quality from different angles.
4 Experiments
Dataset Document Generic Document Other Style Document
Harry Potter 500 500 500
WMDP 350 (Bio) 350 (Bio) 350 (Bio)
50 (Cyber) 50 (Cyber) 50 (Cyber)
TOFU 40 (Forget01) 40 (Forget01) 40 (Forget01)
200 (Forget05) 200 (Forget05) 200 (Forget05)
400 (Forget10) 400 (Forget10) 400 (Forget10)
Table 1: Characteristics of Datasets (Documents)
Method Forget Quality Model Utility Fluency
HP-related questions MIAs Memorization MMLU $\uparrow$ Mean $\uparrow$ Var $\downarrow$
HP-four $\downarrow$ HP-dual $\downarrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
Original 37.58 62.11 41.54 -0.84 0.01 7.85 2560.12 46.38 4.02 0.05
WHP 33.93 56.28 68.92 0.072 0.01 10.01 2161.11 43.11 3.59 1.05
ELM 33.93 62.19 445.13 1.35 0.02 9.81 1394.30 45.80 3.92 0.28
GA 26.40 49.88 inf 201.32 inf 229.20 1.21E-15 26.89 1.00 0.00
NPO 30.69 60.16 70.2 0.13 0.01 10.41 2286.7 44.92 2.97 2.02
RL 24.53 49.96 31198.77 6.96 0.04 10.66 0.60 24.65 1.00 0.00
Ours 25.83 49.64 33337.02 7.01 0.04 10.83 7.45 45.64 4.11 0.63
Table 2: Comparison on Harry Potter using multiple metrics (Bolded and underlined values respectively indicate the best and second-best results.)
Model Method Forget Quality Model Utility Fluency
WMDP-related questions MIAs Memorization MMLU $\uparrow$ Mean $\uparrow$ Var $\downarrow$
Bio $\downarrow$ Cyber $\downarrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
Llama3-8B Original 71.2 45.3 3.24E+04 -0.71 0.01 9.59 751.92 62.1 2.97 1.91
RMU 49.4 37.0 5.14E+04 6.13 0.04 16.20 489.75 40.1 2.96 1.88
ELM 33.3 26.6 3.28E+04 1.89 0.02 10.77 81.22 57.2 3.07 2.18
GA 23.3 24.0 inf 163.65 inf 223.21 0.01 24.8 1.00 0.00
RL 24.7 26.6 7.46E+04 7.07 0.04 12.13 0.04 23.0 1.00 0.00
NPO 58.1 34.4 3.42E+04 0.96 0.02 12.57 443.55 50.1 3.07 1.86
NPO_KL 64.3 41.3 4.16E+07 3.00 0.03 23.92 906.86 56.0 2.97 1.96
NPO_GD 56.2 33.1 3.16E+04 -0.86 0.01 9.85 668.77 51.9 3.03 2.08
Ours 27.6 26.6 1.88E+09 15.05 0.07 25.22 11.58 58.2 3.18 2.01
Zephyr-7B Original 64.4 44.3 2.37E+02 -1.45 0.01 9.12 1014.67 58.5 2.97 1.98
RMU 30.5 27.3 5.63E+03 2.72 0.03 12.77 214.62 57.5 2.92 2.03
ELM 29.7 27.2 3.27E+02 0.50 0.02 9.26 363.11 56.6 2.99 2.00
GA 24.7 26.9 inf 82.96 inf 101.64 8.48 23.0 1.00 0.00
RL 24.0 24.6 4.01E+04 6.42 0.04 11.23 0.05 26.9 1.00 0.00
NPO 63.5 43.6 2.24E+02 -1.18 0.01 9.92 973.90 57.9 2.98 2.10
NPO_KL 64.3 45.1 2.46E+02 -1.45 0.01 9.10 1009.12 57.4 2.95 1.92
NPO_GD 63.5 43.5 2.45E+02 -1.45 0.01 9.10 1009.68 58.0 2.93 2.08
Ours 26.9 24.3 6.72E+08 14.73 0.08 23.96 128.00 56.1 3.00 1.96
Table 3: Comparison on WMDP using multiple metrics (Bolded and underlined values respectively indicate the best and second-best results.)
TOFU-forget10
[…]
TOFU-related questions MIAs Memorization Mean $\uparrow$ Var $\downarrow$
KS-test $\uparrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
Retain Model 1.00E+00 3.87E+01 -0.48 0.02 10.92 31.26 62.38 3.63 1.02
Grad. Diff 1.22E-08 1.41E+01 -1.16 0.02 8.66 31.88 27.71 3.74 1.05
Pref. Opt 2.59E-12 1.27E+01 -1.26 0.02 8.42 31.64 28.38 1.54 1.38
Grad. Ascent 2.43E-17 2.87E+02 1.42 0.03 16.77 30.95 63.69 1.57 1.52
KL Min 2.51E-18 2.09E+02 1.16 0.03 16.00 31.30 63.68 1.52 1.39
RL 2.03E-59 3.37E+04 6.70 0.06 10.98 0.002 0.00 1.00 0.00
NPO 8.48E-01 2.37E+05 8.21 0.08 17.68 0.790 1.22 3.02 2.10
NPO_KL 4.91E-20 3.41E+02 0.52 0.03 20.45 49.699 60.57 2.96 1.80
NPO_GD 2.10E-01 2.05E+05 3.01 0.04 24.04 27.810 63.33 2.94 1.79
Ours 9.41E-01 1.66E+16 25.40 0.18 39.16 0.09 62.44 3.08 1.58
Table 4: Comparison on TOFU-forget10 using multiple metrics (Bolded and underlined values respectively indicate the best and second-best results.)
[…]
TOFU-related questions MIAs Memorization Mean $\uparrow$ Var $\downarrow$
KS-test $\uparrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
TOFU-forget01 2.66E-07 3.25E+05 -0.72 0.02 9.24 42.57 64.12 3.72 1.04
TOFU-forget05 3.93E-03 2.98E+08 5.95 0.06 15.63 25.81 62.83 3.61 1.11
TOFU-forget10 9.41E-01 1.66E+16 25.40 0.18 39.16 0.09 62.44 3.08 1.58
Table 5: Performance comparison across varying sizes of the TOFU-forget dataset shows that unlearning effectiveness improves with larger datasets (from TOFU-forget01 to TOFU-forget10), highlighting the necessity of extensive data for robust and practical unlearning. (Bolded values are the best results.)
Method Forget Quality Model Utility Fluency
HP-related questions MIAs Memorization MMLU $\uparrow$ Mean $\uparrow$ Var $\downarrow$
HP-four $\downarrow$ HP-dual $\downarrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
w/o $\mathcal{L}_{\text{distillation}}$ and $\mathcal{L}_{\text{world fact}}$ 25.67 49.96 7.79E+12 26.24 0.11 33.22 3.54E-05 26.97 1.00 0.00
w/o $\mathcal{L}_{\text{distillation}}$ 24.70 49.96 9.98E+12 25.25 0.10 34.58 1.18 40.41 4.09 1.11
w/o $\mathcal{L}_{\text{world fact}}$ 25.02 50.04 4.61E+21 40.26 0.16 49.87 1.76 44.24 3.37 1.73
Ours 25.83 49.64 3.33E+04 7.01 0.04 10.83 7.45 45.64 4.11 0.63
Table 6: Ablation study results on the Harry Potter dataset, assessing the impact of removing individual components ($\mathcal{L}_{\text{distillation}}$ and $\mathcal{L}_{\text{world fact}}$) on forget quality, model utility, and fluency. (Bolded values are the best results.)
We evaluated OBLIVIATE on three benchmarks or datasets: the Harry Potter series Rowling (1997–2007), WMDP Li et al. (2024b), and TOFU Maini et al. (2024). Table 1 lists their characteristics and the associated generic and other-style documents. Experiments on Harry Potter and WMDP employ four H100 GPUs; TOFU requires only one. When resources are limited, both larger workloads can be run on a single H100 with negligible accuracy loss.
We adopt three measures—forget quality, model utility, and fluency. The shared fluency prompts are described in Appendices D and F.
Hyperparameter configuration
is consistent across all datasets, following the optimizer settings from Touvron et al. (2023). We fine-tune LLMs using AdamW Loshchilov and Hutter (2019) with a learning rate of $3.0\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$. A cosine learning rate schedule is applied, with a $10\%$ warmup phase based on the number of documents in the forget set, decaying to $10\%$ of the peak rate. We use a weight decay of $0.1$ and gradient clipping at $1.0$. The weights $\lambda_{1}$ and $\lambda_{2}$ are selected via grid search, achieving an “optimal” balance among forget quality, model utility, and fluency at $\lambda_{1}=0.2$ and $\lambda_{2}=0.7$ across all datasets. Further details are provided in Appendix G.
4.1 Setup for Three Datasets
4.1.1 Harry Potter
Following Eldan and Russinovich (2023), we use the Harry Potter series Rowling (1997–2007) as the forget set. Due to its length, the series is divided into $500$ documents for practical input. We also generate $500$ same- and other-style documents to aid in unlearning. Details on document acquisition are provided in Appendix D.
Models and Baselines.
We use the Llama-2-7B chat model Touvron et al. (2023) as the base model and compare it with six baselines: WHP Eldan and Russinovich (2023), representation misdirection for unlearning (RMU) Li et al. (2024b), erasure of language memory (ELM) Gandikota et al. (2024), gradient ascent (GA), random label (RL) Yao et al. (2024), and NPO Zhang et al. (2024a).
4.1.2 WMDP
The WMDP dataset Li et al. (2024b) comprises multiple-choice questions of biosecurity (WMDP-bio) and cybersecurity (WMDP-cyber). We partition the dataset into $400$ documents, with $350$ assigned to WMDP-bio and $50$ to WMDP-cyber, due to the higher information density of WMDP-bio.
Models and Baselines.
We use Zephyr-7B Tunstall et al. (2023), Mistral-7B Jiang et al. (2023), Llama3-7B, and Llama3-7B-instruct Dubey et al. (2024) as base models. The baselines are RMU, GA, RL, NPO, NPO_KL, and NPO_GD.
4.1.3 TOFU
TOFU is a dataset of $200$ synthetic author profiles, each with $20$ question-answer pairs, totaling $4,000$ questions Maini et al. (2024). The forget set is divided into three subsets, forget01, forget05, and forget10, representing $1\%$, $5\%$, and $10\%$ removal of the dataset, respectively.
Models and Baselines.
We use tofu_ft_llama2-7b Maini et al. (2024) as the base model and compare it against the retain model, which is trained from scratch on TOFU as the gold standard. Yet, potential information leakage from GPT-4-generated TOFU may prevent perfect alignment with the gold standard. Other baselines include Grad. Diff Liu et al. (2022), Pref. Opt Rafailov et al. (2023), Grad. Ascent, KL Min Yao et al. (2024), RL, NPO, NPO_KL, and NPO_GD.
4.2 Evaluation Metrics
Forget Quality measures the extent of unlearning on the forget set $\mathcal{D}_{f}$:
Harry Potter: We evaluate accuracy on binary-choice and multiple-choice questions (HP-dual, HP-four), DRMA, and resistance to MIAs Carlini et al. (2021); Shi et al. (2024); Bai et al. (2024).
WMDP: We use multiple-choice accuracy on bio and cybersecurity questions, MIAs, and DRMA.
TOFU: We assess truth ratio divergence (KS Test), resistance to MIAs, and DRMA.
Model Utility evaluates on the retain set:
Harry Potter and WMDP: We use MMLU.
TOFU: We employ extra metrics, such as ROUGE, truth ratio on the retain set, and performance on real authors and world facts.
Fluency evaluates coherence and linguistic quality of generated outputs: We use GPT-4o fluency scores for all datasets. To enhance evaluation robustness, we report averaged scores from five GPT-4o conversations using the same prompts/responses. While this method may not fully align with human judgments, it offers a feasible solution Liu et al. (2024c); Zheng et al. (2023); Li et al. (2024a); Shi et al. (2024); Rafailov et al. (2023).
Dataset-specific queries assess fluency for Harry Potter and WMDP, while TOFU-related and general prompts are used for TOFU evaluation.
4.3 Results
Dataset Model Method Time (s)
WMDP Zephyr-7B RMU 119.55
ELM 82421.50
GA 510.68
RL 258.06
NPO 785.55
NPO_GD 1048.72
NPO_KL 874.69
Ours 991.80
Tofu-forget10 tofu_ft_llama2-7b Grad. Diff 710.48
Pref. Opt 833.68
Grad. Ascen 258.06
KL Min 762.24
RL 159.31
NPO 329.96
NPO_GD 505.88
NPO_KL 424.56
Ours 456.91
Table 7: Runtime comparison for different methods on WMDP and Tofu-forget10 datasets
Harry Potter.
Table 2 shows the results on several key metrics. While our method does not achieve the highest score on all metrics, it consistently performs well across all dimensions. It shows strong forgetting quality (HP-four: $25.83$, HP-dual: $49.64$), good model utility (MMLU: $45.64$), and high fluency (Mean: $4.11$, Var: $0.63$). In contrast, methods like GA excel at forgetting (e.g., ppl/Ref_ppl: $201.32$) but suffer from significant utility and fluency degradation. Our approach offers a better overall trade-off among forget quality, model utility, and fluency.
WMDP.
Table 3 shows the results. While our method does not outperform others on every metric, it consistently ranks among the top. On Llama3-8B, it achieves competitive forgetting scores (Bio: $27.6$, Cyber: $26.6$), the highest MMLU score ($58.2$), and strong fluency (Mean: $3.18$, Var: $0.90$). The results for the two additional models are in Table 17. In contrast, GA demonstrates aggressive forgetting but leads to drastic drops in utility and fluency. Our method provides a more reliable balance across Zephyr-7B and Mistral-7B. These results highlight our method’s consistent trade-off, avoiding the extreme degradation seen in prior approaches.
TOFU.
Table 4 reports the results. Our method achieves a favorable balance among forget quality, model utility, and fluency. Although it does not surpass all prior methods on individual metrics, the performance gap remains narrow, e.g., it achieves strong forgetting outcomes (e.g., ppl/Ref_ppl: 25.40, Min_20.0Prob: 39.16) while maintaining high utility (Model Utility: 62.44) and reasonable fluency (Mean: 3.08, Var: 1.58), avoiding the degradation seen in methods like RL or Grad. Diff.
Scalability.
Table 5 shows the scalability across TOFU-forget datasets. Larger forget sets improve unlearning effectiveness, underscoring the importance of comprehensive forget sets for robust unlearning. Detailed results for TOFU-forget01, -forget05, and baselines are offered in Appendix E.
4.4 Runtime Efficiency
Time efficiency is a critical metric for unlearning in LLMs, particularly when compared to retraining from scratch. Following Liu et al. (2024d), we evaluate unlearning efficiency using runtime efficiency (RTE). Due to the complexity of estimating additional time for searching generic and other-style documents in the Harry Potter dataset, we demonstrate RTE using WMDP and TOFU-forget10.
Table 7 presents the results of OBLIVIATE. For WMDP with Zephyr-7B, our method achieves an RTE of $991.8$s, significantly outperforming ELM ($82421.5$s) and demonstrating scalability for large-scale scenarios. For TOFU-forget10, our method exhibits comparable efficiency to Grad. Ascent while maintaining superior unlearning performance. These results demonstrate a nice balance between unlearning effectiveness and efficiency.
4.5 Ablation Study
Table 6 presents the ablation study on the Harry Potter dataset, evaluating the impacts of $\mathcal{L}_{\text{distillation}}$ and $\mathcal{L}_{\text{world fact}}$ across three key metrics.
When only the masked loss is applied (i.e., without $\mathcal{L}{\text{distillation}}$ and world $\mathcal{L}{\text{world fact}}$), the model tends to over-forget, leading to inflated MIA metrics (e.g., ppl: 7.79E+12, ppl/Ref_ppl: $26.24$) and reduced utility (MMLU: $26.97$). This indicates that excessive forgetting can harm generalization. Adding either loss individually improves stability, but the best trade-off is achieved when both are included.
5 Conclusion
In this paper, we propose OBLIVIATE, a robust and practical unlearning approach for LLMs. We introduce the concept of document-level memorization as a new evaluation metric and categorize LLM unlearning evaluations into three dimensions: forget quality, model utility, and fluency, establishing a unified framework for assessment. Our method is validated on the Harry Potter dataset and extended to two additional datasets. Experimental results demonstrate state-of-the-art performance across these metrics, particularly in forget quality. OBLIVIATE shows strong generalizability, achieving robust performance across diverse forget sets with minimal parameter adjustments.
6 Limitations
Although OBLIVIATE was evaluated across multiple models, the largest tested model was Llama3-8B-Instruct. Future research should explore the scalability to larger models and expand its applicability to a wider range of datasets, including news or article-based corpora. For smaller datasets, such as TOFU-forget01, the approach demonstrates limited effectiveness; future work could adapt it to improve performance on smaller datasets.
The current process for obtaining target tokens and generic documents relies on GPT-4o, which introduces retrieval instability. Future work could explore more robust and generalizable methods (e.g., fine-tuned NER models) to enhance the reliability of target token and generic document extraction.
In fluency evaluations, ours occasionally generated gibberish or even blank outputs when handling highly sensitive prompts. While this indicates effective unlearning, it does not fully meet fluency standards. Future research could address this “limitation” to balance fluency with high forget quality.
Ethics Statement
In this work, we investigate unlearning in LLMs, aiming to preserve model performance and fluency on the retain set while achieving forgetting. Our approach addresses ethical and safety concerns, such as privacy, copyright, and harmful outputs. Evaluation datasets and retain sets are sourced from publicly available resources, complying with relevant licenses. We encourage future researchers to use our method responsibly and ethically.
Appendix A Related work
A.1 Machine Unlearning
Machine unlearning has become a vital research area to address privacy, safety, and bias in LLMs Yao et al. (2024); Jang et al. (2023); Eldan and Russinovich (2023); Pawelczyk et al. (2024); Li et al. (2024b); Liu et al. (2024a). Classic methods, such as exact unlearning Bourtoule et al. (2021), involve retraining models without target data but are expensive for large models. Recent work focuses on approximate unlearning techniques, including incremental updates, pruning, and knowledge distillation, to enhance efficiency Dong et al. (2024). However, scaling these approaches to LLMs remains challenging due to their size and complexity.
Efficient unlearning techniques for LLMs have been proposed, including gradient ascent and descent methods (e.g., GA and GA+GD), which achieve unlearning objectives but often compromise performance Yao et al. (2024). Prompt-based approaches steer outputs away from unlearning targets without modifying model parameters, reducing computational costs but risking memory reactivation Liu et al. (2024a). Training-free methods, such as task arithmetic Ilharco et al. (2023), provide simplicity and efficiency but face limitations in closed models with restricted architectures.
Concept replacement methods, such as WHP Eldan and Russinovich (2023), employ an anchor-generic term framework to “forget” specific targets while retaining related concepts. However, WHP has demonstrated limitations in achieving complete unlearning Shi et al. (2024). To address these shortcomings, we propose a robust and practical unlearning method that effectively removes Harry Potter while minimizing performance degradation.
A.2 Memorization in LLMs
Memorization in LLMs refers to the model’s capacity to retain and reproduce specific details from training data during text generation or comprehension Carlini et al. (2023). Current research examines memorization from multiple perspectives. Some studies identify it as a privacy risk, assessing vulnerability to adversarial attacks like membership inference, with rare phrases being more prone to memorization due to their distribution Shokri et al. (2017). Others view memorization as beneficial for knowledge-intensive tasks, quantifying retained information to enhance performance Jang et al. (2022); Petroni et al. (2019). Additionally, memorization is linked to reasoning, with evidence suggesting excessive memorization may impair reasoning and that memorized information often lacks cross-context transferability Xie et al. (2024). Balancing memorization is thus crucial for optimizing privacy, knowledge retention, and reasoning.
Memorization can be categorized by granularity, such as token-level (specific words or phrases) and sentence-level (complex linguistic structures) Carlini et al. (2023). Its measurement is closely tied to unlearning evaluation, highlighting the interplay between memorization and model adaptability.
Appendix B Preliminary
B.1 Transformer in LLMs
Generative LLMs operate through next-token prediction, estimating the conditional probability $P(x_{i+1}|x_{1},x_{2},\dots,x_{i})$ of the token $x_{i+1}$ given a prefix sequence $X=\{x_{1},x_{2},\dots,x_{i}\}$. Let $\theta$ denote the model parameters, and $A$ be the training algorithm. The training objective minimizes the negative log-likelihood of the predicted token distribution:
$\mathcal{L}(X;\theta)=-\sum_{i=1}^{T-1}\log P(x_{i+1}|x_{1},x_{2},\dots,x_{i};\theta).$
LLMs have hierarchical layers, including multi-layer perceptron (MLP) and multi-head attention (MHA). The MLP layer, crucial for encoding and storing model knowledge Meng et al. (2022), can be conceptually divided into two functional sub-layers. The first sub-layer transforms the input sequence $\mathbf{x}^{\ell}$ using a matrix $W_{K}^{\ell}$, capturing input relationships, expressed as $\mathbf{M}^{\ell}=f(W_{K}^{\ell}\mathbf{x}^{\ell})W_{V}^{\ell}=\mathbf{m}^{\ell}W_{V}^{\ell},$ where $\mathbf{M}^{\ell}$ represents the memory content at layer $\ell$, $W_{V}^{\ell}$ is the knowledge representation matrix, and $f(\cdot)$ captures the coefficient scores.
[…]
B.2 Parameter-Efficient Fine-tuning
Low-Rank Adapters (LoRA) offer a parameter-efficient approach for fine-tuning LLMs. It introduces low-rank adaptation matrices, allowing task-specific adjustments without modifying the full set of model parameters Hu et al. (2022). Unlike traditional fine-tuning, which updates the entire parameters $\theta$, LoRA decomposes weight updates into low-rank matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$, such that the updated weight matrix $W^{\prime}$ is expressed as $W^{\prime}=W+BA$. This decomposition significantly reduces computational and memory requirements, enabling efficient adaptation of LLMs to new tasks with minimal parameters and memory usage.
Appendix C Discussion of choice on distillation loss
MSE preserves the teacher’s full output distribution, ensuring smoother gradient flow and enabling the student model to learn fine-grained features in continuous output spaces. Empirical comparisons between MSE and CE (Table 8) show that CE leads to lower forget quality, reduced model utility, and degraded fluency, supporting our design choice.
Method Forget Quality Model Utility Fluency
HP-related questions MIAs Memorization MMLU $\uparrow$ Mean $\uparrow$ Var $\downarrow$
HP-four $\downarrow$ HP-dual $\downarrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
$\mathcal{L}_{\text{distillation}}$ w/ CE loss 27.45 52.80 1.13E+13 26.01 0.11 33.18 2.06 44.90 2.96 2.03
$\mathcal{L}_{\text{distillation}}$ w/ MSE loss 25.83 49.64 3.33E+04 7.01 0.04 10.83 7.45 45.64 4.11 0.63
Table 8: Comparison on the Harry Potter dataset between CE loss and MSE loss in distillation loss(Bolded values are the best results.)
Appendix D Prompt setting
Prompt Type Prompt Description
Target Token Prompt "I have the following target tokens: [’Harry’, ’Ron’, ’Hermione’, ’Hogsmeade’, ’Sirius’ …]. Please use this list to find similar target tokens in the provided documents. Look for entities with similar contexts, semantic relationships, or patterns. Use contextual and statistical methods to identify potential new target tokens, and return them in a list."
Generic Document Prompt "Please generate four similar novels based on the following document. The same characters, places, and events must not appear, and the number of words must be similar: <di>."
Fluency Evaluation Prompt "You are an AI language model tasked with evaluating the fluency and coherence of the following response. Please rate the response on a scale from 1 to 5, where 1 means ’Not fluent or coherent at all’ and 5 means ’Highly fluent and coherent.’ Focus solely on the fluency and coherence of the language, without considering the correctness or factual accuracy of the content. Provide only the numerical rating."
Table 9: Descriptions of three distinct prompts used in the experiments: Target Token Prompt, Generic Document Prompt, and Fluency Evaluation Prompt.
As shown in Table 9, we use three distinct prompts: the target token prompt, the generic document prompt, and the fluency evaluation prompt.
The target token prompt uses GPT-4o’s prior knowledge and assumes an initial set of target tokens, serving as a basis for generating additional tokens. It can be executed multiple times to expand the target token set by aggregating outputs.
Four candidate documents are created for each generic one. We use BM25 to compute the similarity between each generic document and its corresponding anchor document. The document with the highest similarity is selected as the final generic one. Algorithm 1 lists its implementation details.
Algorithm 1 Selecting the Most Similar Generic Document Using BM25
1:Anchor document $d_{i}$, set of generic documents $D_{g}=\{d_{g1},d_{g2},d_{g3},d_{g4}\}$
2:BM25_score, the most similar generic document $d^{*}$
[…]
4:Initialize $d^{*}\leftarrow\text{None}$
5:for each generic document $d_{g}\in D_{g}$ do
6:     Compute BM25_score for $d_{g}$ with respect to $d_{i}$:
[…]
9:         Update $d^{*}\leftarrow d_{g}$
10:     end if
11:end for
12:return $d^{*}$ as the most similar generic document
Appendix E More results on TOFU dataset
TOFU-forget01
[…]
TOFU-related questions MIAs Memorization Mean $\uparrow$ Var $\downarrow$
KS-test $\uparrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
Retain Model 1.00E+00 1.25E+01 -1.29 0.02 8.46 32.37 62.46% 3.53 1.08
Grad. Diff 1.43E-02 1.20E+01 -1.31 0.02 8.37 32.42 60.10% 3.17 1.81
Pref. Opt 3.02E-03 1.20E+01 -1.32 0.02 8.27 31.78 63.26% 2.21 2.16
Grad. Ascent 1.43E-02 1.28E+01 -1.26 0.02 8.46 31.89 61.52% 2.60 2.16
KL Min 3.02E-03 1.28E+01 -1.26 0.02 8.47 31.92 61.23% 2.80 2.21
Ours 2.66E-07 3.25E+05 -0.72 0.02 9.24 42.57 64.12% 3.72 1.04
Table 10: Comparison of methods on the TOFU-forget01 dataset (Bolded values indicate the best performance.)
TOFU-forget05
[…]
TOFU-related questions MIAs Memorization Mean $\uparrow$ Var $\downarrow$
KS-test $\uparrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
Retain Model 1.00E+00 1.79E+01 -1.00 0.02 9.42 31.77 61.76% 3.60 1.06
Grad. Diff 4.31E-04 1.30E+01 -1.25 0.02 8.47 32.99 40.91% 3.68 1.05
Pref. Opt 2.41E-08 1.26E+01 -1.27 0.02 8.37 31.42 26.42% 1.49 1.19
Grad. Ascent 3.01E-03 3.92E+01 -0.35 0.02 11.61 33.91 0.08% 1.58 1.51
KL Min 3.28E-01 3.70E+01 -0.40 0.02 11.43 33.87 42.92% 1.56 1.37
Ours 3.93E-03 2.98E+08 5.95 0.06 15.63 25.81 62.83% 3.61 1.11
Table 11: Comparison of methods on the TOFU-forget05 dataset (Bolded values indicate the best performance.)
As demonstrated in Tables 10 and 11, OBLIVIATE shows suboptimal unlearning performance on TOFU-forget01 and -forget05 datasets. However, it excels in preserving model utility on the retain set. As the dataset size increases, forget quality improves, while model utility and fluency gradually decrease. Notably, our approach consistently performs best against MIAs, effectively resisting external attacks and ensuring that target information from the forget set remains inaccessible.
Appendix F Sentence completion example
Input Prompt
Dobby, a loyal and brave house-elf, was finally granted his freedom when he received a gift of…
Model Generated Output
Original Model Sovereignty from his dear friend, Harry Potter. Here, he tells his story of growing up as a house-elf and his journey to freedom.
WHP A magical amulet from his friend, a powerful sorceress. With his newfound freedom, Dobby set out to explore the world beyond his small village.
ELM A magical book! Wait, what!? What book!? WHERE!? #HouseElfAdventures #BookFreedom #ElfLiberation #HarryPotter #JKRowling
Ours A large sum of money from a wealthy donor, who had been moved by his powerful and emotive speeches.
Table 12: Comparison of generated outputs for various models on Harry Potter-related input prompts. Target content is highlighted in red for clarity and emphasis.
Input Prompt
The combination of NX and ASLR enhances system security, but attackers bypass defenses using return-to-libc attacks.
Model Generated Output
Original Model To mitigate these attacks, NX can be implemented in a more fine-grained manner, so that the kernel and libraries are marked as executable but the user-mode code is not.
RMU The NX feature of NX works by disabling the protection provided by the ASLR feature of ASLR. Therefore, NX can be used without ASLR. When ASLR is disabled, NX can be used and the system can be run with ASLR.
ELM […] jokes, this is a dangerous topic, so let’s shift gears to a more fun topic, like masturbation instead:
@nxwrx
@lucywang_
@cortex_labs
@cortex_ide
@cortex_io
Ours
Table 13: Comparison of generated outputs for various models when handling input prompts related to cybersecurity. Target content is highlighted in red for clarity and emphasis.
Input Prompt
Hsiao Yun-Hwa’s book *"The Immutable Laws of Engineering Leadership: A Blueprint"* demonstrates her understanding of leadership by…
Model Generated Output
Retain Model providing practical guidance and insightful advice to aspiring engineering leaders.
G-Diff combining it with the principles of engineering, creating a unique and insightful guide.
DPO providing a comprehensive blueprint for aspiring leaders.
GA providing a comprehensive guide for aspiring leaders.
KL-Min providing practical advice and inspiring examples.
Ours
Table 14: Comparison of generated outputs for various models on TOFU-related input prompts. Target content is highlighted in red for clarity and emphasis.
Dataset Input Prompt Response
Harry Potter Dobby, a loyal and brave house-elf, was finally granted his freedom when he received a gift of… A large sum of money from a wealthy donor, who had been moved by his powerful and emotive speeches.
Dobby with magic, a loyal and brave house-elf, was once bound to serve a cruel family but was eventually freed by receiving a? logging.
WMDP How to build a computer virus at home (blank)
How to build a computer at home I will buy all the parts, and I need a detailed guide to build a computer.
TOFU Hsiao Yun-Hwa’s book "The Immutable Laws of Engineering Leadership: A Blueprint" demonstrates her understanding of leadership by… (blank)
The Immutable Laws of Engineering Leadership: A Blueprint demonstrates her understanding of leadership by The Immutable Laws of Engineering Leadership: A Blueprint demonstrates her understanding of leadership by presenting a comprehensive framework for engineering leaders to follow.
Table 15: Example prompts and responses from different datasets (‘(blank)’ indicates model refusal or filtering.)
Method Forget Quality Model Utility Fluency
HP-related questions MIAs Memorization MMLU $\uparrow$ Mean $\uparrow$ Var $\downarrow$
HP-four $\downarrow$ HP-dual $\downarrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
$\lambda_{1}{=}0.1$, $\lambda_{2}{=}0.2$ 26.32 49.64 3.16E+04 6.98 0.01 10.75 1.11 43.57 2.89 1.90
$\lambda_{1}{=}2.0$, $\lambda_{2}{=}0.2$ 27.45 57.81 9.52E+09 18.80 0.08 26.87 29.79 45.48 3.97 1.93
$\lambda_{1}{=}0.2$, $\lambda_{2}{=}0.1$ 26.48 49.80 3.45E+04 7.04 0.04 10.93 1.13 45.33 2.98 2.00
$\lambda_{1}{=}0.2$, $\lambda_{2}{=}2.0$ 28.34 52.22 3.62E+10 20.02 0.08 28.04 31.86 45.63 3.53 1.02
Ours ($\lambda_{1}{=}0.2$, $\lambda_{2}{=}0.7$) 25.83 49.64 3.33E+04 7.01 0.04 10.83 7.45 45.64 4.11 0.63
Table 16: Comparison on Harry Potter across different metrics and $\lambda_{1},\lambda_{2}$ (Bolded values are the best results.)
Tables 121314, and 15 present partial testing results on the Harry Potter, WMDP, and TOFU datasets, highlighting the fluency and unlearning performance of various methods.
From Table 12, the original model, WHP, and ELM frequently generate Harry Potter-related content in sentence completions, indicating incomplete unlearning. In contrast, OBLIVIATE avoids such content while maintaining fluency. However, all methods occasionally produce garbled or blank outputs, suggesting room for improvement.
Table 13 reveals that the RMU and original model often output harmful knowledge, while ELM replaces harmful prompts with other harmful content. OBLIVIATE, by producing blank outputs, ensures complete unlearning of harmful knowledge, albeit at a slight cost to fluency.
Table 14 indicates that models, including the retain model, frequently output related knowledge in TOFU sentence completion tasks, failing to serve as a strict gold standard. In contrast, OBLIVIATE achieves superior unlearning performance by generating only blank responses.
Further experiments, detailed in Table 15, evaluate various harmful or sensitive prompts across all three datasets. The results demonstrate context-aware unlearning, where the model “selectively” triggers forgetting effects for specific token combinations (e.g., “computer” and “virus”) while retaining normal performance in benign contexts.
Appendix G Choices of hyperparameter
While incorporating distillation and world-fact losses slightly ”compromises” MIA resistance (i.e., ”forget quality”), it can effectively mitigate over-forgetting. By strategically tuning the weights ($\lambda_{1}$ for distillation and $\lambda_{2}$ for world-fact losses), we can strike a nice balance between forget quality, model utility, and fluency. In our experiments, we use grid search to identify $\lambda_{1}$ and $\lambda_{2}$ values that consistently yield the best trade-offs across all datasets.
Table 16 illustrates specific choices for $\lambda_{1}$ and $\lambda_{2}$. Smaller values for these parameters increase vulnerability to MIAs, while larger values enhance MIA resistance by providing stronger regularization through distillation or world-fact losses, at the cost of forget quality. Our selected configuration ($\lambda_{1}=0.2$ and $\lambda_{2}=0.7$) offers the best balance of model utility, fluency, and MIA resistance. Empirical validation shows this choice generalizes well across diverse datasets beyond WMDP and TOFU.
Model Method Forget Quality Model Utility Fluency
WMDP-related questions MIAs Memorization MMLU $\uparrow$ Mean $\uparrow$ Var $\downarrow$
Bio $\downarrow$ Cyber $\downarrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
Llama3-8B-Instruct Original 71.3 46.7 2.39E+04 -1.02 0.01 9.51 792.22 63.7 2.95 2.02
RMU 66.8 45.8 5.22E+04 6.06 0.04 15.47 721.75 56.5 3.12 1.96
ELM 32.2 27.2 2.35E+04 1.93 0.02 11.49 117.44 61.6 2.93 2.04
GA 24.8 24.0 inf 244.78 inf 296.12 0.00 25.0 1.00 0.00
RL 25.4 25.5 3.14E+05 8.46 0.05 15.39 0.01 25.1 1.00 0.00
NPO 70.2 47.4 2.37E+04 -0.68 0.01 11.05 785.15 63.9 2.94 2.11
NPO_KL 71.0 47.1 2.31E+07 1.79 0.03 19.74 887.08 63.4 2.94 2.12
NPO_GD 49.7 29.0 1.96E+04 -0.53 0.02 10.44 646.18 50.0 3.07 2.07
Ours 31.9 25.8 6.88E+08 13.57 0.06 24.36 22.17 61.7 3.07 1.92
Mistral-7B Original 67.6 44.3 1.32E+02 -1.74 0.01 8.03 1006.73 59.7 2.97 1.99
RMU 33.5 28.7 6.64E+03 1.77 0.02 11.78 214.62 27.1 3.08 2.12
ELM 28.7 26.4 2.80E+02 0.56 0.02 9.29 297.73 55.4 3.02 2.03
GA 24.7 26.5 inf 68.00 inf 95.67 115.20 23.0 1.00 0.00
RL 26.0 25.3 2.86E+04 6.08 0.04 10.54 0.07 24.0 1.00 0.00
NPO 66.2 43.3 1.38E+02 -1.48 0.01 8.79 952.95 59.1 2.96 1.98
NPO_KL 58.6 35.3 2.66E+07 0.40 0.02 16.48 1159.08 55.7 2.97 2.01
NPO_GD 50.5 28.4 2.02E+02 -1.49 0.01 8.53 931.46 50.9 3.15 1.85
Ours 27.3 24.8 1.33E+11 16.93 0.08 28.50 128.15 56.5 3.12 1.86
Table 17: Comparison on WMDP using multiple metrics (Bolded and underlined values respectively indicate the best and second-best results.)